# LoongServe: Efficiently Serving Long-context Large Language Models with Elastic Sequence Parallelism

### v1

@@ 9. Conclusion
To serve long-context LLM under the dynamic workload, we propose elastic sequence parallelism and build LoongServe, which provides a set of elastic scaling mechanisms without additional overhead and a scalable scheduling algorithm for ESP. Evaluation across diverse datasets shows that compared to SOTA solutions, LoongServe significantly improves the total throughput, input token throughput, and output token latency simultaneously.
Appendix A Other Optimizations
Relaxed FCFS Dispatching. Although FCFS is simple and efficient enough for the short-context LLMs, it is not always optimal for the long-context LLMs. For requests with short input lengths, even if they can execute, previous requests acquiring more resources may block them, leading to long queuing time. For requests with long input lengths, even if they can execute, there may be few elastic instances available for them, leading to long execution time.
[…]
Greedy Execution. In each iteration, different batches may have different execution times. For example, the decoding phase is typically much faster than the prefill phase. If these independent batches need to wait for the slowest one, there will be much idle time. In this case, the global manager allows faster batches to continue executing more iterations called greedy iterations, as long as the slowest batch does not finish. If a faster batch finishes, the global manager will schedule pending requests to fill the idle time as long as they finish before that of the slowest normal iteration. As a result, it does not cause starvation.

### v2

@@ 9. Conclusion
To serve long-context LLMs under dynamic workloads, we propose ESP and build LoongServe, which provides a set of elastic scaling mechanisms without additional overhead and a scalable scheduling algorithm for ESP. Evaluation across diverse real-world datasets shows that compared to existing solutions, LoongServe significantly improves the performance of the prefill phase and decoding phase simultaneously.
Acknowledgments. We sincerely thank our shepherd, Rong Chen, and the anonymous reviewers for their valuable feedback. This work was supported by the National Key Research and Development Program of China under the grant number 2022YFB4500700, National Natural Science Foundation of China under the grant numbers 62172008 and 62325201. Xin Jin is the corresponding author. Bingyang Wu, Shengyu Liu, Yinmin Zhong, Xuanzhe Liu and Xin Jin are also with the Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education.